OpenStudy (anonymous):

WILL MEDAL!!!!!! Solve the problem. The formula d = rt gives the distance traveled in time t at rate r. If a bicyclist rides for 2.3 hours and travels 41 miles, what was her average speed?

OpenStudy (anonymous):

A. 9.43 miles per hour B. 5.6 miles per hour C. 17.83 miles per hour D. 38.7 miles per hour

OpenStudy (anonymous):

i think its c....

OpenStudy (anonymous):

@ageta

OpenStudy (ageta):

mean you cant go 38.7 mph on a bike or 17.83 mph

OpenStudy (anonymous):

so not d?

OpenStudy (ageta):

distance = time x rate... which is d = rt if we want to find the rate .. we can write it as rate = distance / time which is r = d/t where d = distance = 41 miles t = time = 2.3 hours so... r= 41/2.3 = 17.83 miles per hour

OpenStudy (ageta):

C is the answer

OpenStudy (anonymous):

oh ok thx alot

OpenStudy (anonymous):

you're the best @ageta

OpenStudy (ageta):

tnxs

OpenStudy (anonymous):

can you help me with another?

OpenStudy (ageta):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The formula P = ns gives the formula for the perimeter of a regular polygon with n sides and side length s. What is the length of each side of a regular octagon with a perimeter of 54.4 cm? A regular octagon has 8 equal sides.

OpenStudy (anonymous):

Any idea of what it could be?

OpenStudy (ageta):

just want to make sure its right i got 435.2

OpenStudy (anonymous):

wow thx alot!

OpenStudy (anonymous):

in cm right?

OpenStudy (ageta):

on problem

OpenStudy (anonymous):

its in cm's right?

OpenStudy (ageta):

yes

OpenStudy (anonymous):

ok good thx

OpenStudy (anonymous):

YES I GOT AN 80% THX SOOOOOOOOOOO MUCH!